# Multiple solutions to a semilinear elliptic equation with a sharp change of sign in the nonlinearity

### v1

@@ Abstract
We consider a nonautonomous semilinear elliptic problem where the power nonlinearity is multiplied by a discontinuous coefficient that equals one inside a bounded open set $\Omega$ and it equals minus one in its complement. In the slightly subcritical regime, we prove the existence of concentrating positive and nodal solutions. Moreover, depending on the geometry of $\Omega$, we establish multiplicity of positive solutions. Finally, in the critical case, we show the existence of a blow-up positive solution when $\Omega$ has nontrivial topology. Our proofs rely on a Lyapunov-Schmidt reduction strategy which in these problems turns out to be remarkably simple. We take this opportunity to highlight certain aspects of the method that are often overlooked and present it in a more accessible and detailed manner for nonexperts.
Key words and phrases. Sign-changing nonlinearity, slightly subcritical exponent, critical exponent.
@@ 1 Introduction
where $12<q\leq p:=\frac{n+2}{n-2}=2^{*}-$ 1, $\Omega\subset\mathbb{R}^{n}$ is a bounded (not necessarily connected) smooth open subset, $n\geq 3$, and   $Q_{\Omega}:={\mathds{1}}_{\Omega}-{\mathds{1}}_{\mathbb{R}^{n}\setminus\Omega}$,  i.e.,
@@
$-\Delta u=u^{p},\qquad u\in D^{1,2}(\mathbb{R}^{n}),$ (1.2)
@@
[…]
@@ 2.1 A decomposition of the space
$-\Delta z=pU_{\delta,\xi}^{p-1}z,\qquad z\in D^{1,2}(\mathbb{R}^{n}).$ (2.2)
@@
$\Pi_{\delta,\xi}:D^{1,2}(\mathbb{R}^{n})\to K_{\delta,\xi}\quad\text{ and }\quad\Pi_{\delta,\xi}^{\perp}:D^{1,2}(\mathbb{R}^{n})\to K_{\delta,\xi}^{\perp}.$
@@
$(i):$  Since $s>n/(n-2)$, a direct calculation shows that $Z_{\delta,\xi}^{i}\in L^{s}(\mathbb{R}^{n})$ for all $i=0,1,\ldots,n$.
@@ 3.1 Existence of the error
the statement follows from Lemmas 2.1 and  2.2. ∎
@@
is well defined. Applying the mean value theorem to the function $t\mapsto f_{\varepsilon}(U_{\delta,\xi}+t\phi_{1}+(1-t)\phi_{2})$ and arguing as in the proof of Lemma 3.1$(i)$ we get
@@
This shows that  $T^{\prime}_{\delta,\xi}(\phi_{\varepsilon}^{d,\xi})$  tends to 00 as $\varepsilon\to 0$ in the Banach space space of continuous linear maps $K^{\perp}_{\delta,\xi}\cap L^{s}(\mathbb{R}^{n})\to K^{\perp}_{\delta,\xi}\cap L^{s}(\mathbb{R}^{n})$. Since the set of linear homeomorphisms is open in this space (see [12, (8.3.2)]), we have that $\mathrm{id}_{K_{\delta,\xi}^{\perp}\cap L^{s}(\mathbb{R}^{n})}-T^{\prime}_{\delta,\xi}(\phi_{\varepsilon}^{d,\xi})$ is a linear homeomorphism for $\varepsilon$ sufficiently small. Hence, $\partial_{\psi}F_{\varepsilon}(d,\xi,P_{\delta,\xi}^{-1}\phi_{\varepsilon}^{d,\xi})$ is a linear homeomorphism for $\varepsilon$ sufficiently small, as claimed. Now we may apply the implicit function theorem [12, (10.2.1)] to derive that $(d,\xi)\mapsto P_{d,\xi}^{-1}\phi_{\delta,\xi}$ is of class $\mathcal{C}^{1}$. It follows that $(d,\xi)\mapsto\phi_{\delta,\xi}$ is of class $\mathcal{C}^{1}$. This completes the proof. ∎
@@ 3.2 Finding the right parameters
To prove (3.16) recall that $K_{\delta,\xi}=\mathrm{span}\{Z^{i}_{\delta,\xi}:i=0,1,\ldots,n\}$. Note that $\langle Z^{i}_{\delta,\xi},Z^{j}_{\delta,\xi}\rangle=0$ if $i\neq j$, and that $\|Z^{i}_{\delta,\xi}\|=\|Z^{i}_{1,0}\|$ does not depend on $\delta$ and $\xi$. As $Z^{i}_{\delta,\xi}$ solves (2.2), we have that $-\Delta(\partial_{\xi_{j}}Z^{i}_{\delta,\xi})=p\partial_{\xi_{j}}(U_{\delta,\xi}^{p-1}Z^{i}_{\delta,\xi})$. Testing this equation with $\partial_{\xi_{j}}Z^{i}_{\delta,\xi}$ yields that $\|\partial_{\xi_{j}}Z^{i}_{\delta,\xi}\|=O(\delta^{-1})$. Now,
@@
Note that $c_{q,\Omega}$ is the least energy of a nontrivial solution to the Dirichlet problem $-\Delta u=|u|^{q-1}u$,  $u\in D^{1,2}_{0}(\Omega)$. It is well known that $\lim_{q\to p}c_{q,\Omega}=\frac{1}{n}S^{\frac{n}{2}}$; see [5, Lemma 4.1]. Therefore, $\limsup\limits_{p\to q}c_{q}\leq\frac{1}{n}S^{\frac{n}{2}}$. This proves $(i)$.
$(ii):$  A well known argument shows that, if $u$ is a nontrivial solution to (2.10) and $J_{q}(u)<2c_{q}$, then $u$ does not change sign; see, for instance, [5, Proof of Theorem A]. Assume that $u\geq 0$. By standard elliptic regularity (see [7, Lemma 3.3]), $u\in W^{2,r}_{loc}(\mathbb{R}^{n})\cap\mathcal{C}^{1,\alpha}_{loc}(\mathbb{R}^{n})$ for all $r\in[1,\infty)$ and $\alpha\in(0,1)$. Then, the strong maximum principle for strong solutions [14, Theorem 9.6] applied to the equation $-\Delta u+{\mathds{1}}_{\mathbb{R}^{n}\setminus\Omega}u^{q}={\mathds{1}}_{\Omega}u^{q}$ yields that $u>0$ in $\mathbb{R}^{n}$. ∎
@@
for some function $r$ that depends only on $\varepsilon$. The Ansatz managed to join the summand $-\varepsilon(c_{2}+c_{3}\ln\delta)$ and $c_{4}\delta^{n}\psi_{\Omega}(\xi)$ to produce the function $c_{4}d^{\frac{1}{n}}\psi_{\Omega}(\xi)-c_{3}\ln d$ that has a global minimum (in $d$).
@@ 4 The slightly subcritical problem: existence of nodal solutions
We build a solution (3.1) with the Ansatz $U_{\delta_{1},\xi_{1}}-U_{\delta_{1},\xi_{1}}+\phi,$ namely, two bubbles with opposite signs plus a small error. Considering the leading terms in the expansion of the reduced energy (see Proposition 4.3), we make the following choices for the variables involved in the Ansatz. We assume that the concentration parameters $\delta_{1}$ and $\delta_{2}$ satisfy that
@@
$\xi_{j}=\eta_{j}+\tau_{j}\nu_{j}\in\Omega,\ \hbox{where}\ \eta_{j}\in\partial\Omega,\ \nu_{i}\ \hbox{is the inner normal at $\eta_{j}$,}$ (4.3)
$\tau_{j}=t_{j}\tau,\ \hbox{with}\ \tau:=\varepsilon^{\frac{2}{(n-2)(n+1)}}\ \hbox{and}\ t_{j}\in(0,\infty).$
@@
$\Pi:D^{1,2}(\mathbb{R}^{n})\to K\quad\text{ and }\quad\Pi^{\perp}:D^{1,2}(\mathbb{R}^{n})\to K^{\perp}.$
@@
[…]
@@
[…]
@@
$\int_{\mathbb{R}^{n}}|\nabla(U_{1}-U_{2})|^{2}=\int_{\mathbb{R}^{n}}|\nabla U_{1}|^{2}+|\nabla U_{1}|^{2}-2\nabla U_{1}\nabla U_{2}=\int_{\mathbb{R}^{n}}|\nabla U_{1}|^{2}+|\nabla U_{1}|^{2}-2U_{1}^{p}\nabla U_{2}.$
@@
$+2\frac{1}{p+1-\varepsilon}\int_{\mathbb{R}^{n}\setminus\Omega}|U_{2}-U_{2}|^{p+1-\varepsilon}-2\frac{1}{p+1}\int_{\mathbb{R}^{n}\setminus\Omega}|U_{1}-U_{2}|^{p+1}$
@@
$I_{4}$ $:=2\frac{1}{p+1-\varepsilon}\int_{\mathbb{R}^{n}\setminus\Omega}|U_{2}-U_{2}|^{p+1-\varepsilon}-2\frac{1}{p+1}\int_{\mathbb{R}^{n}\setminus\Omega}|U_{1}-U_{2}|^{p+1}.$
@@
$I_{4}$ $=2\frac{1}{p+1-\varepsilon}\left(\int_{\mathbb{R}^{n}\setminus\Omega}|U_{2}-U_{2}|^{p+1-\varepsilon}-\int_{\mathbb{R}^{n}\setminus\Omega}|U_{1}-U_{2}|^{p+1}\right)+2\left(\frac{1}{p+1-\varepsilon}-\frac{1}{p+1}\right)\int_{\mathbb{R}^{n}\setminus\Omega}|U_{1}-U_{2}|^{p+1}$
@@
[…]
@@ 5 The critical problem in a domain with a shrinking hole
In the following we use $C>0$ to denote possibly different constants independent of $\varrho$. Setting $\varepsilon=0$ in the proof of Lemma 3.1$(i)$ we see that the new nonlinear operator $\mathscr{N}$ defined above satisfies
@@
where $B_{\varrho}$ is the ball of radius $\varrho$ centered at 00. The rest of the proof is like that of Proposition 3.3. ∎
@@
where $B_{\varrho}$ is the ball of radius $\varrho$ centered at 00. A straightforward computation shows that
@@
The point $(d_{0},0)$ with $d_{0}=(\frac{\mathfrak{b}_{2}}{\mathfrak{b}_{1}})^{\frac{1}{2n}}$ is the only critical point of the function $\Phi$. It is a nondegenerate saddle point, stable under $\mathcal{C}^{1}$-perturbations. Therefore, for small enough $\varrho$, the reduced energy $\widetilde{J}_{\varrho}$ has a critical point that converges to $(d_{0},0)$ as $\varrho\to 0$, and problem (1.10) has a solution $u_{\varrho}$ concentrating at 00, whose energy satisfies $J_{\varrho}(u_{\varrho})\to\mathfrak{a}=\frac{1}{n}S^{\frac{n}{2}}$. By Lemma 3.6$(ii)$, we have that $u_{\varrho}$ is strictly positive in $\mathbb{R}^{n}$. ∎

### v2

@@ Abstract
We consider a nonautonomous semilinear elliptic problem where the power-type nonlinearity is multiplied by a discontinuous coefficient that takes the value one inside a bounded open set $\Omega$ and minus one in its complement. In the slightly subcritical regime, we prove the existence of concentrating positive and nodal solutions. Moreover, depending on the geometry of $\Omega$, we establish multiplicity of positive solutions. Finally, in the critical case, we show the existence of a blow-up positive solution when $\Omega$ has nontrivial topology. Our proofs rely on a Lyapunov-Schmidt reduction strategy which in these problems turns out to be remarkably simple. We take this opportunity to highlight certain aspects of the method that are often overlooked and present it in a more accessible and detailed manner for nonexperts.
Key words and phrases. Nonautonomous nonlinearity, critical and slightly subcritical problem, blow up, Lyapunov–Schmidt.
@@ 1 Introduction
where $1<q\leq p:=\frac{n+2}{n-2}=2^{*}-1$, $\Omega\subset\mathbb{R}^{n}$ is a bounded (not necessarily connected) smooth open subset, $n\geq 3$, and   $Q_{\Omega}:={\mathds{1}}_{\Omega}-{\mathds{1}}_{\mathbb{R}^{n}\setminus\Omega}$,  i.e.,
@@
$\Delta u=u^{p},\qquad u\in D^{1,2}(\mathbb{R}^{n}),$ (1.2)
@@
Our first result shows that the slightly subcritical problem has a positive solution for any bounded open set $\Omega\subset\mathbb{R}^{n}$ and characterizes its concentration behavior as $\varepsilon\to 0$. We write $\|\cdot\|$ for the norm in $D^{1,2}(\mathbb{R}^{n})$ and $|\cdot|_{q}$ for the norm in $L^{q}(\mathbb{R}^{n})$.
@@ 2.1 A decomposition of the space
$\Delta z=pU_{\delta,\xi}^{p-1}z,\qquad z\in D^{1,2}(\mathbb{R}^{n}).$ (2.2)
@@
[…]
@@
$(i):$  Since s>n/(n2)s>n/(n-2), a direct calculation shows that $Z_{\delta,\xi}^{i}\in L^{s}(\mathbb{R}^{n})$ for all $i=0,1,\ldots,n$.
@@ 3.1 Existence of the error
the statement follows from Lemmas 2.1 and  $2.$ ∎
@@
is well defined. Applying the mean value theorem to the function $t\mapsto f_{\varepsilon}(U_{\delta,\xi}+t\phi_{1}+(1-t)\phi_{2})$ and arguing as in the proof of Lemma 3.1(i)(i) we get
@@
This shows that  $T^{\prime}_{\delta,\xi}(\phi_{\varepsilon}^{d,\xi})$  tends to 0 as $\varepsilon\to 0$ in the Banach space space of continuous linear maps $K^{\perp}_{\delta,\xi}\cap L^{s}(\mathbb{R}^{n})\to K^{\perp}_{\delta,\xi}\cap L^{s}(\mathbb{R}^{n})$. Since the set of linear homeomorphisms is open in this space (see [12, (8.3.2)]), we have that $\mathrm{id}_{K_{\delta,\xi}^{\perp}\cap L^{s}(\mathbb{R}^{n})}-T^{\prime}_{\delta,\xi}(\phi_{\varepsilon}^{d,\xi})$ is a linear homeomorphism for $\varepsilon$ sufficiently small. Hence, $\partial_{\psi}F_{\varepsilon}(d,\xi,P_{\delta,\xi}^{-1}\phi_{\varepsilon}^{d,\xi})$ is a linear homeomorphism for $\varepsilon$ sufficiently small, as claimed. Now we may apply the implicit function theorem [12, (10.2.1)] to derive that $(d,\xi)\mapsto P_{d,\xi}^{-1}\phi_{\delta,\xi}$ is of class $\mathcal{C}^{1}$. It follows that $(d,\xi)\mapsto\phi_{\delta,\xi}$ is of class $\mathcal{C}^{1}$. This completes the proof. ∎
@@ 3.2 Finding the right parameters
To prove (3.16) recall that $K_{\delta,\xi}=\mathrm{span}\{Z^{i}_{\delta,\xi}:i=0,1,\ldots,n\}$. Note that $\langle Z^{i}_{\delta,\xi},Z^{j}_{\delta,\xi}\rangle=0$ if $i\neq j$, and that $\|Z^{i}_{\delta,\xi}\|=\|Z^{i}_{1,0}\|$ does not depend on $\delta$ and $\xi$. As $Z^{i}_{\delta,\xi}$ solves (2.2), we have that $\Delta(\partial_{\xi_{j}}Z^{i}_{\delta,\xi})=p\partial_{\xi_{j}}(U_{\delta,\xi}^{p-1}Z^{i}_{\delta,\xi})$. Testing this equation with $\partial_{\xi_{j}}Z^{i}_{\delta,\xi}$ yields that $\|\partial_{\xi_{j}}Z^{i}_{\delta,\xi}\|=O(\delta^{-1})$. Now,
@@
Note that $c_{q,\Omega}$ is the least energy of a nontrivial solution to the Dirichlet problem $\Delta u=|u|^{q-1}u$,  $u\in D^{1,2}_{0}(\Omega)$. It is well known that $\lim_{q\to p}c_{q,\Omega}=\frac{1}{n}S^{\frac{n}{2}}$; see [5, Lemma 4.1]. Therefore, $\limsup\limits_{p\to q}c_{q}\leq\frac{1}{n}S^{\frac{n}{2}}$. This proves $(i)$.
$(ii):$  A well known argument shows that, if $u$ is a nontrivial solution to (2.10) and $J_{q}(u)<2c_{q}$, then $u$ does not change sign; see, for instance, [5, Proof of Theorem A]. Assume that $u\geq 0$. By standard elliptic regularity (see [7, Lemma 3.3]), $u\in W^{2,r}_{loc}(\mathbb{R}^{n})\cap\mathcal{C}^{1,\alpha}_{loc}(\mathbb{R}^{n})$ for all $r\in[1,\infty)$ and $\alpha\in(0,1)$. Then, the strong maximum principle for strong solutions [14, Theorem 9.6] applied to the equation $\Delta u+{\mathds{1}}_{\mathbb{R}^{n}\setminus\Omega}u^{q}={\mathds{1}}_{\Omega}u^{q}$ yields that $u>0$ in $\mathbb{R}^{n}$. ∎
@@
for some function $r$ that depends only on $\varepsilon$. The Ansatz managed to join the summand $\varepsilon(c_{2}+c_{3}\ln\delta)$ and $c_{4}\delta^{n}\psi_{\Omega}(\xi)$ to produce the function $c_{4}d^{\frac{1}{n}}\psi_{\Omega}(\xi)-c_{3}\ln d$ that has a global minimum (in $d$).
@@ 4 The slightly subcritical problem: existence of nodal solutions
We build a solution (3.1) with the Ansatz $U_{\delta_{1},\xi_{1}}-U_{\delta_{1},\xi_{2}}+\phi,$ namely, two bubbles with opposite signs plus a small error. Considering the leading terms in the expansion of the reduced energy (see Proposition 4.3), we make the following choices for the variables involved in the Ansatz. We assume that the concentration parameters $\delta_{1}$ and $\delta_{2}$ satisfy that
@@
$\xi_{j}=\eta_{j}+\tau_{j}\nu_{j}\in\Omega,\ \hbox{where}\ \eta_{j}\in\partial\Omega,\ \nu_{j}\ \hbox{is the inner normal at $\eta_{j}$,}$ (4.3)
$\tau_{j}=t_{j}\tau,\ \hbox{with}\ \tau=\varepsilon^{\frac{2}{(n-2)(n+1)}}\ \hbox{and}\ t_{j}\in(0,\infty).$
@@
[…]
@@
$\left\{\begin{aligned} \Pi^{\perp}\left\{U_{1}-U_{2}+\phi-i^{*}\left[Q_{\Omega}f_{\varepsilon}\left(U_{1}-U_{2}+\phi\right)\right]\right\}&=0,\\ \Pi\left\{U_{1}-U_{2}+\phi-i^{*}\left[Q_{\Omega}f_{\varepsilon}\left(U_{1}-U_{2}+\phi\right)\right]\right\}&=0.\end{aligned}\right$. (4.6)
@@
$\int_{\mathbb{R}^{n}\backslash\Omega}|U_{i}|^{\alpha\beta}$ $=\int_{\mathbb{R}^{n}\backslash\Omega}\left|\frac{\alpha_{n}\delta_{i}^{{\frac{n-2}{2}}}}{(\delta_{i}^{2}+|x-\eta_{i}-\tau_{i}\nu_{i}|^{2})^{\frac{n-2}{2}}}\right|^{\alpha\beta}\,dx$
$=C\delta_{i}^{{\frac{n-2}{2}}\alpha\beta}\tau_{i}^{n-\alpha\beta(n-2)}\int_{\mathbb{R}^{n}\backslash(\frac{\Omega-\eta_{i}}{\tau_{i}})}\left|\frac{1}{((\frac{\delta_{i}}{\tau_{i}})^{2}+|y-\nu_{i}|^{2})^{\frac{n-2}{2}}}\right|^{\alpha\beta}\,dy$
@@
$\int_{\mathbb{R}^{n}}|\nabla(U_{1}-U_{2})|^{2}=\int_{\mathbb{R}^{n}}|\nabla U_{1}|^{2}+|\nabla U_{1}|^{2}-2\nabla U_{1}\nabla U_{2}=\int_{\mathbb{R}^{n}}|\nabla U_{1}|^{2}+|\nabla U_{1}|^{2}-2U_{1}^{p}U_{2}.$
@@
$+2\frac{1}{p+1-\varepsilon}\int_{\mathbb{R}^{n}\setminus\Omega}|U_{1}-U_{2}|^{p+1-\varepsilon}-2\frac{1}{p+1}\int_{\mathbb{R}^{n}\setminus\Omega}|U_{1}-U_{2}|^{p+1}$
@@
$I_{4}$ $:=2\frac{1}{p+1-\varepsilon}\int_{\mathbb{R}^{n}\setminus\Omega}|U_{1}-U_{2}|^{p+1-\varepsilon}-2\frac{1}{p+1}\int_{\mathbb{R}^{n}\setminus\Omega}|U_{1}-U_{2}|^{p+1}.$
@@
$I_{4}$ $=2\frac{1}{p+1-\varepsilon}\left(\int_{\mathbb{R}^{n}\setminus\Omega}|U_{1}-U_{2}|^{p+1-\varepsilon}-\int_{\mathbb{R}^{n}\setminus\Omega}|U_{1}-U_{2}|^{p+1}\right)+2\left(\frac{1}{p+1-\varepsilon}-\frac{1}{p+1}\right)\int_{\mathbb{R}^{n}\setminus\Omega}|U_{1}-U_{2}|^{p+1}$
@@
$=\int_{\mathbb{R}^{n}}\frac{\alpha_{n}^{\frac{n+2}{n-2}}\delta_{1}^{\frac{n+2}{2}}}{(\delta_{1}^{2}+|x-\xi_{1}|^{2})^{\frac{n+2}{2}}}\frac{\alpha_{n}\delta_{2}^{\frac{n-2}{2}}}{(\delta_{2}^{2}+|x-\xi_{2}|^{2})^{\frac{n-2}{2}}}\,dx+O(\delta^{n})$
$=\alpha_{n}^{\frac{n+2}{n-2}+1}\delta_{1}^{\frac{n-2}{2}}\delta_{2}^{\frac{n-2}{2}}\int_{\mathbb{R}^{n}}\frac{1}{(1+|y|^{2})^{\frac{n+2}{2}}}\frac{1}{(\delta_{2}^{2}+|\delta_{1}y+\xi_{1}-\xi_{2}|^{2})^{\frac{n-2}{2}}}\,dy+O(\delta^{n})$
@@ 5 The critical problem in a domain with a shrinking hole
In the following we use $C>0$ to denote possibly different constants independent of $\varrho$. Setting $\varepsilon=0$ in the proof of Lemma 3.1(i)(i) we see that the new nonlinear operator $\mathscr{N}$ defined above satisfies
@@
where $B_{\varrho}$ is the ball of radius $\varrho$ centered at 0. The rest of the proof is like that of Proposition 3.3. ∎
@@
where $B_{\varrho}$ is the ball of radius $\varrho$ centered at 0. A straightforward computation shows that
@@
The point $(d_{0},0)$ with $d_{0}=(\frac{\mathfrak{b}_{2}}{\mathfrak{b}_{1}})^{\frac{1}{2n}}$ is the only critical point of the function $\Phi$. It is a nondegenerate saddle point, stable under $\mathcal{C}^{1}$-perturbations. Therefore, for small enough $\varrho$, the reduced energy $\widetilde{J}_{\varrho}$ has a critical point that converges to $(d_{0},0)$ as $\varrho\to 0$, and problem (1.10) has a solution $u_{\varrho}$ concentrating at 0, whose energy satisfies $J_{\varrho}(u_{\varrho})\to\mathfrak{a}=\frac{1}{n}S^{\frac{n}{2}}$. By Lemma 3.6(ii)(ii), we have that $u_{\varrho}$ is strictly positive in $\mathbb{R}^{n}$. ∎